Evaluate the piecewise defined function at the indicated values.
f(x)={x(square) if x<0
{x+1 if x_>0
f(-2), f(-1), f(0), f(1),. f(2)

Answer by stanbon(75887)

f(x)={x(square) if x<0
{x+1 if x_>0
--------------------------------
f(-2)= (-2)^2=4
------------------------------
f(-1)= (-1)^2=1
------------------------------
f(0)=0+1=1
------------------------------
f(1)=1+1=2
--------------------------
f(2)=2+1=3
-------------------------
Cheers,
Stan H.